OpenStudy (anonymous):

What do you notice about the vertex and the line of symmetry for this equations y = x2 + 4x + 1?

OpenStudy (cwrw238):

y = (x + 2)^2 -4 + 1 y = (x + 2)^2 - 3

OpenStudy (cwrw238):

from this you can find the vertex and line of symmetry